OpenStudy (anonymous):

x/2y-y/2x --------- 2/y+2/x

OpenStudy (amistre64):

looks like someones trying to get there homework/test done ;)

OpenStudy (amistre64):

how you like the new openstudy sstar?

OpenStudy (amistre64):

multiply top and bottom by 2xy right? x^2-y^2 --------- 4x+4y

OpenStudy (anonymous):

\[\frac{\frac{x}{2y}- \frac{y}{2x}}{\frac{2}{y}+\frac{2}{x}} \]\[= \frac{x^2-y^2}{2yx}.\frac{xy}{2x+2y} = \frac{xy(x-y)(x+y)}{4yx(x+y)} = \frac{(x-y)}{4}\] Correct me if I'm wrong ^_^

OpenStudy (anonymous):

I like it very much amistre ^_^, it's very nice!

OpenStudy (amistre64):

divide by your favorite way of dividing...long hand prolly x/4 -y/4 ----------------- 4x +4y | x^2 -y^2 -x^2 -xy ----------- -xy -y^2 xy + y^2 ---------- 0 x/4 - y/4 = (x-y)/4 yor right :)

OpenStudy (anonymous):

awesome ^_^~

OpenStudy (amistre64):

i think you get a medal for it ;)

OpenStudy (anonymous):

thank you all so much this really helped!!

OpenStudy (anonymous):

most welcome :)